Oxygen (molar mass 32 kg/kmol) expands reversibly in a cylinder behind a piston at a constant pressure of 3.163 bar. The volume initially is 0.0112 mand finally is 0,0323 m3 , the initial temperature is 17.6 0C. Calculate the heat supplied during the expansion in the correct unit to two decimal places. Assume oxygen to be a perfect gas and take the specific heat at constant pressure as = 0.917kJ / kg K and molar gas constant as = 8.314 J / kmol K.


1
Expert's answer
2021-08-24T06:09:17-0400

Solution.

"p \times V = n \times R \times T"

"n = \frac{p \times V}{R \times T}"

"n = \frac{3.163 \times 101325 \times 0.0112}{8.31 \times (17.6+273.14)} = 1.49 \ mol"

"\frac{V1}{T1} = \frac{V2}{T2}"

"T2 = \frac{T1 \times V2}{V1}"

T2 = 838.48 K

"Q = n \times Cp \times \Delta T"

"Q = 1.49 \times 0.917 \times 10^3 \times (838.48-290.74) = 748393.59 \ J = 748.39 \ kJ"

Answer:

Q = 748.39 kJ